# Teola: Towards End-to-End Optimization of LLM-based Applications

### v1

@@ 1. Introduction
Figure 2. Real-world LLM-based application workflows, showcasing typical definition styles in current frameworks (pairag, ; llamaindex, ; langchain, ).
@@
We implement Teola’s prototype primarily using Ray (moritz2018ray, ) for distributed scheduling and execution, with various libraries for the execution engines. We evaluate Teola with diverse datasets and applications, including search engine-empowered generation and document question answering using both naive and advanced RAG. Comprehensive testbed experiments demonstrate that Teola can achieve up to a 2.09x speedup in end-to-end latency compared to existing schemes, including our distributed implementation of Llamaindex (llamaindex, ) with Ray and its advanced version that incorporates module parallelization and enhanced LLM execution.
@@ 2.1. LLM-based Applications
LLM apps are more than just LLM. Despite their great generation capabilities, LLMs are not a panacea. Their training datasets are inevitably not up-to-date, leading to knowledge gaps and hallucination issues (azurerag, ; lewis2020retrieval, ). They also lack abilities to interact directly with the environment, that is they are not directly capable of sending an email though it can draft the email message (wu2023autogen, ; hong2023metagpt, ; shen2023hugginggpt, ). Thus real-world applications often need to integrate additional tools with LLMs to be practically usable.
We show in Figure 2 four typical LLM-based applications (apps). Figure 2(a) demonstrates a search engine-empowered generation app, where the LLM utilizes the search engine to answer questions that are beyond its knowledge scope (bing, ; jeong2024adaptiverag, ; tan2024small, ). It employs a proxy and a judge model to determine if the search engine needs to be called. Figure 2(b) illustrates a generic LLM agent, where the LLM interacts with various tool APIs to execute the plan it formulates (e.g. draft and send emails using the user’s account credentials) (hong2023metagpt, ; wu2023autogen, ). Figures 2(c) and 2(d) showcase document question answering (QA) with naive and advanced RAG, respectively. RAG is arguably the most popular technique to enhance LLM apps with many production uses (pairag, ; azurerag, ). Here the documents uploaded by users are ingested as chunks into the vector database as the domain knowledge base (lewis2020retrieval, ; pairag, ; azurerag, ), after processed by the embedding models. This step is known as indexing. The LLM uses the relevant chunks retrieved from the vector database to generate answers. The advanced version (Figure 2(d)) leverages LLM-based query expansion to refine and broaden new queries (jagerman2023queryexpansion, ; gao2023retrievalsurvey, ), thereby enhancing search accuracy, and subsequently reranks all retrieved chunks for the expanded queries to synthesize a precise final answer. As seen before in 1, the LLM may not be the only performance bottleneck of these complex application pipelines. Carefully orchestrating the various components of the workflow is thus critical.
@@ 2.2. Fine-grained Orchestration of LLM Apps
Many frameworks such as LlamaIndex (llamaindex, ), Langchain (langchain, ), and enterprise solutions such as PAI-RAG (pairag, ) and Azure-RAG (azurerag, ) have emerged to facilitate the creation and orchestration of LLM applications. They naturally adopt module-level orchestration in the sense that each app is defined and scheduled as a simple chain of modules, as depicted in Figure 3. Each module is executed independently with backend engines. Coarse-grained module-level chaining is easy to use, but inherently limited for optimizing the complex workflows for best performance. It overlooks the larger design space of jointly optimizing the modules, especially by exploiting the intricate dependencies among the internal operations of the individual modules.
@@ 3.2. APIs
Workflow template. A workflow template defines the essential components for an app and their execution flow (line 8-20). Developers specify components involved in a workflow, outlining required engines, roles, and input-output configurations. These components can be further annotated with optimization hints, such as batchable (for batched inputs that operate independently) and splittable (for output that can be divided into independent partial outputs). The >>much-greater-than$>$ > operator establishes execution sequence between components, ensuring dataflow correctness. The resulting template serves as the basis for constructing and optimizing a finer-grained graph for queries with different configurations.
@@ 7. Evaluation
Figure 8. End-to-end performance of search engine-empowered generation (top), document QA with naive RAG (middle), and document QA with advanced RAG (bottom). Subtitles for each subfigure indicate the dataset and core LLM.
Testbed setup. We allocate model-based engines (e.g., LLMs) and model-free engines with GPUs and CPU-only resources, respectively. Each engine instance used for embeddings or other non-LLM models are each hosted on a single NVIDIA 3090 24GB GPU. For LLMs, each instance of llama-2-7B and llama-2-13B (touvron2023llama2model, ) is deployed on 1 and 2 NVIDIA 3090 GPUs, respectively. Each instance of llama-30B (touvron2023llamamodel, ) is deployed on 2 NVIDIA A800 80GB GPUs. The network bandwidth between any physical servers is 100 Gbps.
Baseline. To our knowledge, few studies have specifically focused on optimizing LLM-based workflows in distributed settings. Therefore, we compare Teola with the following frameworks based on LlamaIndex (llamaindex, ):
  • LlamaDist: a distributed version of LlamaIndex that we implemented with Ray, defining a chain of task modules to construct an application pipeline. Each task module invokes requests to different distributed backend engines. This implementation integrates Ray with LlamaIndex’s orchestration approach, utilizing the same engines as Teola but differing in the granularity of orchestration.
  • LlamaDistPC (parallel & cache-reuse): an advanced LlamaDist variant that examines the predefined pipeline and manually parallelizes independent modules for concurrent execution. It also incorporates prefix caching for LLM to avoid recomputation for partial instructions in the prompt, as proposed in some previous works (zheng2023sglang, ; liu2024optimizingquery, ; gim2023promptcache, ).
@@
  • Document QA with naive RAG (Figure 2(c)): Users input documents or webpages along with the question. The app segments the documents into chunks (default chunk size: 256, chunk overlap:30), embeds them with the bge-large-en-v1.5 model (xiao2023bgeembedding, ), and stores them in a vector database (postgresql & pgvector). It retrieves the most relevant chunks (default: top 3) to generate a response with a tree-based mode. The workload (i.e. question and documents/webpage) is synthesized from Finqabench (Finqabench, ) and TruthfulQA (lin2021truthfulqa, ) datasets using a Poisson distribution.
  • Document QA with advanced RAG (Figure 2(d)): Extending the second app, a query expansion is used (the core LLM) to rewrite and expand the original query into multiple new queries (default: 3), improving retrieval accuracy. A reranker (bge-reranker-large (xiao2023bgeembedding, )) evaluates the similarity between retrieved chunks, with each query searching for 16 chunks and determining the top 3 overall. These top chunks are fed into the core LLM for generation in a refine mode as mentioned in 4.2.
@@ 7.1. End-to-end Performance
Search engine-empowered generation. Figure 8 (top row) shows Teola outperforming the other four schemes by up to 1.79x. Teola’s efficiency is attributed to parallelizable partial prefilling for instructions and questions for both the judge and core LLM, and effective batching coordination for different engines. In contrast, LlamaDist executes modules sequentially and struggles with request scheduling for multiple queries. PO’s focus on per-invocation latency results in longer queue times under high request rates, while TO generally performs better in these scenarios. LlamaDistPC fails to benefit from parallelization due to the lack of explicit parallelization across modules. Its prefix caching for partial instructions (typically around 60 tokens) provides limited benefit, as prefix caching is most advantageous when prefixes are significantly longer (promptflow, ; zheng2023sglang, ).
@@
Document QA with naive RAG. Figure 8 (middle row) demonstrates that Teola outperforms the other four schemes by up to 1.62x at low request rates and 1.46x at high rates. LlamaDist executes modules sequentially without specific optimizations while LlamaDistPC enables limited parallelization (indexing and query embedding modules) and partial instruction KV cache reuse, performing slightly better than LlamaDist. Regarding scheduling, PO is better than TO at low request rates due to its focus on per-invocation latency, but performance suffers at high rates. Furthermore, the app introduces intricate request relationships. Both the indexing and query embedding modules utilize the embedding model. Meanwhile, the LLM synthesizing module makes three initial requests followed by a subsequent request to construct the tree synthesis. If overlooked, these can cause batching inefficiencies, leading to reduced goodput, similar to TO. Conversely, Teola leverages the e-graph, incorporating pipelining to split compute-heavy tasks like large embeddings for document chunks, while also exploring more parallelization opportunities, such as four partial prefilling. Additionally, Teola’s topology-aware batching captures dependencies and correlations among requests linked to different primitives, facilitating effective batching for each engine.
@@
Document QA with advanced RAG. This app is the most complex in our settings, yet it provides ample opportunities to demonstrate the effectiveness of Teola. It leverages aggressive optimization techniques such as parallelization at different levels (e.g., independent dataflow branches and partial prefilling for different LLM calls) and pipelining (e.g., breaking large embeddings into smaller ones and splitting the decoding process in query expansion into three partial decodes), as shown in Figure 6. In contrast, LlamaDist runs sequentially with a simple run-to-completion paradigm, missing opportunities to reduce end-to-end latency. LlamaDistPC improves parallelization across the indexing and query expansion modules and reuses partial KV cache but still fails to explore the full optimization potential like Teola. Additionally, similar to naive RAG, both LlamaDist and LlamaDistPC struggle to efficiently coordinate requests whether in PO or TO, whereas Teola performs well. Overall, Teola outperforms others by up to 2.09x at low request rates and 1.68x at high request rates, as shown in Figure 8 (bottom row).
@@ 7.2. Ablation Study
Graph optimization. As shown in Figure 9, we compare the performance of Teola under different scenarios, i.e., with or without parallelization (Pass 1 & 3) and pipelining (Pass 2 & 4) optimization as mentioned in 4. The left figure shows the average end-to-end latency of a single query and explicitly demonstrates that both parallelization and pipelining effectively capture optimization opportunities and reduce latency. This holds true as well in a request trace with different request rates.
@@ 7.3. Overhead Analysis
To demonstrate that the overhead incurred by Teola, we provide a breakdown of latency by profiling the different parts in the real critical path of execution. This analysis covers document QA with advanced RAG on the TruthfulQA dataset at varying request rates. It includes latency measurements for graph optimization, communication, queuing, and primitive execution. The results clearly show that the graph optimization overhead is minimal (1.3% ~3%) with optimization cache reuse, and the communication overhead is low ( 3.1% ~6.2%) compared to the total latency. As the request rate increases, more latency is attributed to the queuing time for certain operations. These indicate that Teola incurs negligible overhead.
@@
Dynamic graph. While Teola’s ahead-of-time graph optimization offers benefits, adapting to dynamic patterns like generation with reflection (madaan2024selfreflection, ), iterative retrieval in RAG (liu2024iterativere, ), and agent-determined workflows (kim2023llmcompiler, ; shen2023hugginggpt, ) is challenging due to their unpredictable patterns and the difficulty of capturing the entire primitive-level graph prior to execution.
Coupling with the backends. To enable finer-grained orchestration, we had to modify several engine-side mechanisms, such as supporting decomposed primitive operations and certain batching strategies. These modifications required extra engineering efforts compared to existing frameworks like LlamaIndex (llamaindex, ) and Langchain (langchain, ) that decouple orchestration and execution and work with pluggable engines. However, these additional efforts enhance performance. Besides, at the interface level, Teola hides optimization details from users so as to be user-friendly.
@@
Multi-app co-orchestration. The current design focuses on a single app but has the potential to be extended for co-orchestrating multiple apps sharing common engines within a cluster, such as RAG and LLM dialogue apps. By analyzing their distinct dataflow graphs and requirements, we can achieve broader system-wide optimizations. Optimizing the cross applications’ performance is left as the future work.

### v2

@@ 1. Introduction
(e) Contextual retrieval from Anthropic (contextual-retrieval, )
Figure 2. Common real-world LLM-based application workflows obtained from orchestration frameworks (pairag, ; llamaindex, ; langchain, ) and Anthropic (contextual-retrieval, ).
@@
We implement Teola’s prototype primarily using Ray (moritz2018ray, ) for distributed scheduling and execution, along with various libraries for the execution engines. We evaluate Teola using diverse datasets and applications, including search engine-empowered generation, document question answering using both naive and advanced RAG, and a newly released production workflow from Anthropic (contextual-retrieval, ). Comprehensive testbed experiments demonstrate that Teola can achieve up to a 2.09x speedup in end-to-end latency compared to existing schemes. These schemes include our distributed implementation of Llamaindex (llamaindex, ) with Ray and its advanced version that incorporates module parallelization and enhanced LLM execution, as well as another agentic framework, AutoGen (wu2023autogen, ).
@@ 2.1. LLM-based Applications
LLM apps are more than just LLM. Despite their great generation capabilities, LLMs are not a panacea. Their training datasets are inevitably not up-to-date, leading to knowledge gaps and hallucination issues (azurerag, ; lewis2020retrieval, ). They also lack abilities to interact directly with the environment, that is they are not directly capable of sending an email though it can draft one (wu2023autogen, ; hong2023metagpt, ; shen2023hugginggpt, ). Thus real-world applications often need to integrate additional tools with LLMs to be practically usable.
Figure 2 illustrates five LLM-based application workflows, four of which are widely used in enterprise-production and open-source projects, and one recently released production workflow from Anthropic. Figure 2(a) demonstrates a search engine-empowered generation app, where the LLM utilizes the search engine to answer questions that are beyond its knowledge scope (bing, ; jeong2024adaptiverag, ; tan2024small, ). It employs a proxy and a judge model to determine if the search engine needs to be called. Figure 2(b) illustrates a generic LLM agent, where the LLM interacts with various tool APIs to execute the formulated plan (e.g. draft and send emails using the user’s account credentials) (hong2023metagpt, ; wu2023autogen, ). Figures 2(c) and 2(d) showcase document question answering (QA) with naive and advanced RAG, respectively. RAG is arguably the most popular technique to enhance LLM apps with many production uses (pairag, ; azurerag, ). Here the documents uploaded by users are ingested as chunks into the vector database as the domain knowledge base (lewis2020retrieval, ; pairag, ; azurerag, ), after processed by the embedding models. This step is known as indexing. The LLM uses the relevant chunks retrieved from the vector database to generate answers. The advanced version (Figure 2(d)) leverages LLM-based query expansion to refine and broaden new queries (jagerman2023queryexpansion, ; gao2023retrievalsurvey, ), thereby enhancing search accuracy, and subsequently reranks all retrieved chunks for the expanded queries to synthesize a precise final answer. Lastly, Figure 2(e) presents Anthropic’s contextual retrieval workflow (contextual-retrieval, ). The LLM generates contextual summaries for each document chunk using additional content from the document. These summaries are prepended to the chunks, ensuring that each fragment maintains coherence with the original document when processed independently.
Workflow Count Proportion
SE generation 16 53.33%
LLM agent 13 43.33%
Doc QA w/ naive RAG 26 86.67%
Doc QA w/ advanced RAG 23 76.67%
Table 1. The occurrence count and proportion of the four workflows (Figure 2(a)–(d)) in 30 sampled GitHub projects with more than 1K stars.
We also conduct a small case study to validate the prevalence of the first four workflows by searching GitHub for open-source projects using the keyword “LLM applications.” We select 30 best-matched projects with more than 1,000 stars, including notable ones such as Haystack (haystack, ), AutoGen (wu2023autogen, ), and Langchain (langchain, ). We manually check these projects to see if the four workflows or their similarities were present; it is common for them to contain multiple workflows. The results, presented in Table 1, illustrate the popularity of these LLM-based workflows. We observe that most workflows occur in more than half of the sampled projects, with RAG-based QA being the most popular in current open-source projects.
With the increasing adoption of such workflows, it is evident that the LLM may not be the sole performance bottleneck in complex application pipelines, as discussed earlier in 1. Therefore, carefully orchestrating the various components of the workflow is critical.
@@ 2.2. Fine-grained Orchestration of LLM Apps
Several frameworks, such as LlamaIndex (llamaindex, ), Langchain (langchain, ), and enterprise solutions like PAI-RAG (pairag, ) and Azure-RAG (azurerag, ), have emerged to facilitate the creation and orchestration of LLM applications. They naturally adopt module-level orchestration, where each app is defined and executed as a simple chain of modules, as depicted in Figure 3. Each module is executed independently on their backend engines. While coarse-grained module-level chaining is easy to implement and use, it is inherently limited in optimizing complex workflows for performance. It overlooks the larger design space of jointly optimizing the modules, particularly by exploiting the intricate dependencies among the internal operations of the individual modules. Moreover, several agentic-like frameworks (wu2023autogen, ; autogpt, ) feature a more compact and coarse-grained pattern, in which multiple agents with distinct roles working on distinct sub-tasks are connected with a predefined workflow to complete a complex task. These frameworks have not explored opportunities of end-to-end optimization across agents or sub-tasks either.
@@ 3.2. APIs
Workflow definition. While Teola abstracts away the complexity of graph optimization and scheduling, developers must first define a high-level workflow template with runtime-rendered placeholders as input for each application, similar to existing frameworks (langchain, ; llamaindex, ; promptflow, ; wu2023autogen, ). Using Teola’s APIs (lines 8-20), developers specify essential components with their required engines, roles, and input-output configurations. These components can be annotated with optimization hints, such as batchable (for independent inputs) and splittable (for divisible outputs). The >>much-greater-than$>$ > operator establishes execution sequences across components to ensure dataflow correctness. This workflow template serves as the foundation for constructing and optimizing detailed execution graphs that adapt to various query configurations.
@@ 5.2. Engine Scheduler
5.3. Co-located Applications
Teola is designed to optimize individual app workflows. When applications are co-located in the same infrastructure (zhang2023shepherd, ; gujarati2020clockwork, ), they are orchestrated by Teola independently. That is, the e-graphs of different applications containing different or identical primitives with distinct app-level and query-level metadata, are treated by Teola in the same manner as in single-app scenarios. This leaves out the potential for further optimization when for example some co-located applications share the LLM engines and the model so advanced techniques like KV cache sharing across different applications may further improve throughput. We plan to explore these opportunities in future work. For completeness, we evaluate Teola’s capability to handle queries in co-located app scenarios with a discussion in 7.2.
@@ 7. Evaluation
Figure 8. End-to-end performance of search engine-empowered generation (1st row), document QA with naive RAG (2nd row), document QA with advanced RAG (3rd row) and contextual retrieval (4th row). Subcaption of each subfigure indicates the dataset and core LLM.
Testbed setup. We allocate model-based engines (e.g., LLMs) and model-free engines with GPUs and CPU-only resources, respectively. Each engine instance used for embeddings or other non-LLM models are each hosted on a single NVIDIA 3090 24GB GPU. For LLMs, each instance of gemma-2-2B, llama-2-7B, and llama-2-13B (touvron2023llama2model, ) is deployed on 1, 1, and 2 NVIDIA 3090 GPUs, respectively. Each instance of llama-30B (touvron2023llamamodel, ) is deployed on 2 NVIDIA A800 80GB GPUs. The network bandwidth between any physical servers is 100 Gbps.
Baseline. To our knowledge, few studies have specifically focused on optimizing LLM-based workflows execution in distributed settings. Therefore, we compare Teola with the following popular frameworks and their variants:
  • LlamaDist: a Ray-based distributed implementation of LlamaIndex (a popular framework with over 36k GitHub stars) that defines a chain of task modules to construct an application pipeline. Each task module invokes requests to different distributed backend engines. This implementation integrates Ray with LlamaIndex’s orchestration approach, utilizing the same engines as Teola but differing in the granularity of orchestration.
  • LlamaDistPC (parallel & cache-reuse): an advanced LlamaDist variant that examines the predefined pipeline and manually parallelizes independent modules for concurrent execution. It also incorporates prefix caching for LLM to avoid recomputation for partial instructions in the prompt, as proposed in some previous works (zheng2023sglang, ; liu2024optimizingquery, ; gim2023promptcache, ; lin2024parrot, ).
  • AutoGen: a popular framework for LLM applications with over 32k stars on GitHub. It constructs various agents for an application, with each agent managing several system modules (e.g., a retrieval agent handling indexing, query embedding, and searching) and communicating with each other following a pre-defined graph.
@@
  • Document QA with naive RAG (Figure 2(c)): Users input documents or webpages along with the question. The app segments the documents into chunks (default size: 256, overlap:30), embeds them with the bge-large-en-v1.5 model (xiao2023bgeembedding, ), and stores them in a vector database (postgresql & pgvector). It retrieves the most relevant chunks (default: top 3) to generate a response with a tree-based mode. The workload (i.e. question and documents/webpage) is synthesized from Finqabench (Finqabench, ) and TruthfulQA (lin2021truthfulqa, ) datasets using a Poisson distribution.
  • Document QA with advanced RAG (Figure 2(d)): Extending the second app, a query expansion is used (core LLM) to rewrite and expand the original query into multiple new queries (default: 3), improving retrieval accuracy. A reranker (bge-reranker-large (xiao2023bgeembedding, )) evaluates the similarity between retrieved chunks, with each query searching for 16 chunks and determining the top 3 overall. These top chunks are fed into the core LLM for generation in a refine mode as mentioned in 4.2.
  • Contextual retrieval (Figure 2(e)): Extending the naive RAG, this workflow incorporates contextualization using a lightweight LLM, gemma-2-2B (team2024gemma, ), to process each chunk before indexing. To optimize time and cost, each chunk is contextualized with its four neighboring chunks, generating a concise context to prepend to each chunk. A reranker, identical to the one in the advanced RAG, reranks the 32 fetched chunks. The top 3 chunks are then input into the core LLM for one-shot generation.
@@ 7.1. End-to-end Performance
Search engine-empowered generation. Figure 8 (1st row) shows Teola outperforming the other six schemes by up to 1.79x. Teola’s efficiency is attributed to parallelizable partial prefilling for instructions and questions for both the judge and core LLM, and effective batching coordination for different engines. In contrast, LlamaDist executes modules sequentially and struggles with request scheduling for multiple queries. PO’s focus on per-invocation latency results in longer queue times under high request rates, while TO generally performs better in these scenarios. LlamaDistPC fails to benefit from parallelization due to the lack of explicit parallelization across modules. Its prefix caching for partial instructions (typically around 60 tokens) provides limited benefit, as prefix caching is most advantageous when prefixes are significantly longer (promptflow, ; zheng2023sglang, ). AutoGen defines agents for the proxy model, judge model, search engine, and LLM synthesizer in this case, resulting in a structure similar to LlamaDist and therefore causing similar performance.
@@
Document QA with naive RAG. Figure 8 (2rd row) demonstrates that Teola outperforms the other six schemes by up to 1.62x at low request rates and 1.67x at high rates. LlamaDist executes modules sequentially without specific optimizations, while LlamaDistPC enables limited parallelization (indexing and query embedding modules) and partial instruction KV cache reuse, performing slightly better than LlamaDist. While AutoGen’s workflow, consisting of a retrieval agent and synthesizer agent, performs adequately at low request volumes, its simplified orchestration architecture results in extended queuing delays when handling high request loads. Regarding scheduling, PO outperforms TO at low request rates due to its focus on per-invocation latency, but its performance suffers at high rates. Furthermore, the app introduces intricate request relationships, with both the indexing and query embedding modules utilizing the embedding model, and the LLM synthesizing module making three initial requests followed by a subsequent request to construct the tree synthesis. If overlooked, these relationships can cause batching inefficiencies and reduced goodput, similar to TO. In contrast, Teola leverages the e-graph, incorporating pipelining to split compute-heavy tasks like large embeddings for document chunks, while also exploring more parallelization opportunities, such as four partial prefilling. Additionally, Teola’s topology-aware batching captures dependencies and correlations among requests linked to different primitives, facilitating effective batching for each engine.
@@
Document QA with advanced RAG. This app is the most complex in our settings, yet it provides ample opportunities to demonstrate the effectiveness of Teola. It leverages aggressive optimization techniques such as parallelization at different levels (e.g., independent dataflow branches and partial prefilling for different LLM calls) and pipelining (e.g., breaking large embeddings into smaller ones and splitting the decoding process in query expansion into three partial decodes), as shown in Figure 6. In contrast, LlamaDist runs sequentially with a simple run-to-completion paradigm, missing opportunities to reduce end-to-end latency. LlamaDistPC improves parallelization across the indexing and query expansion modules and reuses partial KV cache but still fails to explore the full optimization potential like Teola. For AutoGen, it would define four agents for retrieval, reranking, query expansion, and LLM synthesizer in this workflow, but it suffers from high request load due to its inability to pipeline and parallelize operations effectively. Additionally, similar to naive RAG, all other baselines struggle to efficiently coordinate requests whether in PO or TO, whereas Teola performs well. Overall, Teola outperforms others by up to 2.09x at low request rates and 2.03x at high request rates, as shown in Figure 8 (3rd row).
Contextual retrieval. This workflow mainly adds contextualization with a lightweight LLM for chunks before indexing and a reranking step after searching, compared to naive RAG. However, due to the massive volume of chunks to process, the contextualization occupies a significant portion of time for all schemes. Although little graph-level optimization could be applied to the contextualization part, Teola’s application-aware scheduling and graph optimization on other components still help it achieve 1.06x-1.59x speedup compared to other baselines, as shown in Figure 8 (4th row), which further validates the applicability of Teola.
7.2. Co-located Applications
Figure 9. The average latency for two apps under a co-location scenario using llama-2-13B as the core LLM on truthfulQA dataset.
Beyond evaluating Teola’s performance in single-application scenarios, we also assess its capability to handle concurrent queries across multiple applications. To demonstrate this, we present a scenario where two document QA applications—a naive RAG-based one and an advanced RAG-based one—operate simultaneously on the same underlying infrastructure. These apps share similar backend engines but differ in their calling semantics, with the advanced RAG-based QA requiring an additional reranking engine.
We compare the average latency of Teola against the stronger baseline, LlamaDistPC, at a request rate of 3 requests per second per application. Figure 9 illustrates the average latency for queries from both applications, revealing that Teola maintains a performance advantage over LlamaDistPC in multi-app co-location scenarios. Teola achieves a 1.2x to 1.55x latency speedup across the two apps, highlighting the consistency of its capabilities in handling multiple application queries. This can be attributed to Teola’s app-agnostic graph optimization and app-aware scheduling techniques.
Furthermore, in real-world deployments, applications could have varying priorities. For example, LLM chat services prioritize shorter latency to enhance user experience, while document summarization services may tolerate longer latencies. These differing priorities can influence the order of request processing within the infrastructure. The varying objectives among applications can be encoded as metadata within their generated primitives and used to facilitate priority-based scheduling or resource-constrained scheduling in each engine scheduler. However, a detailed exploration of this aspect is beyond the scope of this paper.
@@ 7.3. Ablation Study
Graph optimization. As shown in Figure 10, we compare the performance of Teola under different scenarios, i.e., with or without parallelization (Pass 1 & 3) and pipelining (Pass 2 & 4) optimization as in 4. The left figure shows the average latency of a single query and explicitly demonstrates that both parallelization and pipelining effectively capture optimization opportunities and reduce latency. This holds true as well in a request trace with different request rates.
@@ 7.4. Overhead Analysis
Partial Prefilling Full Prefilling Total Single Prefilling
76.03 (200) 215.89 (800) 291.92 (1000) 260.36 (1000)
217.67 (850) 222.66 (850) 440.33 (1700) 414.09 (1700)
582.95 (2500) 159.65 (500) 742.60 (3000) 720.15 (3000)
Table 3. Execution efficiency comparison between decomposed prefilling (partial and full prefilling) of Teola (left) and single complete prefilling (right) for different input sizes with llama-2-7B. Values represent execution time in millionsecond, with input size in parentheses.
We have shown that Teola provides substantial performance improvements over existing solutions. Here in order to better understand Teola’s overheads we provide a more detailed analysis in mainly three aspects: graph optimization, data communication, and execution efficiency. First, e-graph optimization incurs new overhead for each query, but this is often minor due to the limited number of primitives in an LLM-based workflow and the ability to cache the results for queries of the same app. Second, communication across primitives also leads to higher compared to existing coarse-grained approach. Teola’s pre-scheduling mechanism mitigates this, and the pipelined execution also allows partially conveyed data to be processed earlier. Third, decomposing certain operations, such as LLM prefilling, may compromise execution efficiency (e.g., repeatedly moving KV cache into SRAM or failing to fully utilize the GPU) compared to a single complete operation. Nevertheless, this performance sacrifice is minimal, as demonstrated by a comparison in Table 3, where the slowdown is only between 3.11% and 12.12%. Moreover, the gained parallelism further outweighs the slight loss and provides a larger end-to-end speedup.
We provide a latency breakdown for document QA with advanced RAG on the TruthfulQA dataset as a case study. We observe that graph optimization overhead is minimal, ranging from 1.3% to 3% of the total latency, when leveraging caching optimizations. Similarly, the communication overhead remains low, accounting for only 3.1% to 6.2%. Under increased request rates on fixed resources, queuing latency dominates end-to-end time, further diminishing the relative impact of these overheads on overall execution time.
@@
In this section, we point out some limitations of Teola and identify future research directions.
Dynamic workflows. LLM-based workflows typically require upfront human design but offer stable performance, leading to widespread adoption in both open-source projects (llamaindex, ; langchain, ; haystack, ; lazyllm, ) and production environments (alibaba_llmapp_observation, ; pairag, ; azurerag, ). Teola is primarily designed for such structured scenarios. However, autonomous agents operating without human intervention have emerged, wherein the LLM plans and executes actions dynamically without a fixed workflow. Teola’s ahead-of-time graph optimization would struggle with dynamic workflows, especially those involving self-reflection (madaan2024selfreflection, ), as it cannot capture the complete primitive-level graph before execution. Adapting Teola for such dynamic workflows is a very promising avenue for future research.
Coupling with the backends. To enable finer-grained orchestration, we have to modify several engine-side mechanisms (e.g. support decomposed primitive operations and certain batching strategies), which requires extra engineering efforts compared to frameworks (llamaindex, ; langchain, ) that decouple the orchestration and execution and work with pluggable engines. This is an essential trade-off between performance and modularity. These optimization details remain encapsulated from the user’s perspective.